OpenStudy (tiffanymak1996):

lim(x->0) xsin(1/x) use the squeezing theorem

Parth (parthkohli):

Always, \(-1 \le\sin \le 1\). Does that give you a proper hint? :)

OpenStudy (tiffanymak1996):

i don't quite understand...

Parth (parthkohli):

Have you ever heard of the Squeeze Theorem?(It's also known as 'Sandwich' or 'Pinching' Theorem).

OpenStudy (tiffanymak1996):

yeah, that's the hint given in the question.

Parth (parthkohli):

So you haven't heard of it. Hmm. I recommend you to see this link. http://tutorial.math.lamar.edu/Classes/CalcI/ComputingLimits.aspx

OpenStudy (tiffanymak1996):

wait, i think i know what you mean.

Parth (parthkohli):

Great!

Parth (parthkohli):

Can you make an inequality to start with?

OpenStudy (lgbasallote):

im still wondering how @ParthKohli knows the squeeze

Parth (parthkohli):

Also to remember: \(-1 \le \sin x \le 1 \Longrightarrow {-1 \le \sin{1 \over x} \le 1}\)

Parth (parthkohli):

@lgbasallote I started Calculus a week ago.

Parth (parthkohli):

Heh - no need to worry! This is not a copied answer!

OpenStudy (lgbasallote):

but this is not "just calculus"

Parth (parthkohli):

This is a part of limits, and I am studying Calculus on Paul's Notes.

OpenStudy (tiffanymak1996):

the answer's 0 right?

OpenStudy (tiffanymak1996):

what about lim(x->0+) (lnx)/(cotx)